OpenStudy (ashley1nonly):

The football is thrown when the player was D distance away. The ball is thrown at a 45 degree angle. What are the equation for the runner and the ball thrown. How hard does the ball need to be thrown in order to be caught. How far will the ball be thrown?

OpenStudy (ashley1nonly):

or you can look at picture on this website the player ran with a constant velocity http://aplusphysics.com/community/index.php/blog/168/entry-1016-physics-of-throwing-a-football/

OpenStudy (anonymous):

hello there :) for what parameter of the runner you want an eqn?

OpenStudy (anonymous):

lets imagine that the initial velocity of the runner is u and final velocity (the velocity when he reaches to get the ball) be v okay now the distance is given as D.. while starting off at velocity u he accelerates at an acceleration of a to reach final velocity v now D=u*t + 1/2(a*t^2) where t is the time

OpenStudy (ashley1nonly):

ay= - g ax=0 vy= -gt + v0 vy=v0 xy= -1/2gt^2+v0t+x0 xx=v0t y component= -1/2gt^2+v0sin(45)t x component= v0cos(45)t this is what i got for my kinematics equations. is it right

OpenStudy (ashley1nonly):

that was just for the football player throwing the ball for the runner i got a= a v=at+vo x=1/2at^2+v0t+x0

OpenStudy (anonymous):

yes you are right :) its a projectile motio for the ball so u need to consider the x and y and by the way for the runner make v0=0 cos hes at rest

OpenStudy (ashley1nonly):

so his x component will be just x=t

OpenStudy (ashley1nonly):

ok it should be v=at

OpenStudy (anonymous):

for the runner it is v=at s=1/2gt^2 a=a

OpenStudy (anonymous):

yup :)

OpenStudy (ashley1nonly):

i have all of these equations, now how do i solve for my answer with how hard the ball needs to be thrown and at what distance will it be caught

OpenStudy (ashley1nonly):

i know time relates them all

OpenStudy (anonymous):

okay.. now hard does the ball need to thrown is the first qn.. here we have to relate both the distances.. because inorder for the ball to be caught the distance covered by the ball should be equal to the distance covered by the runner

OpenStudy (ashley1nonly):

so the distance would have to be D=d+w because the ball was thrown when he was at a distance d and he continued to run a distance of w

OpenStudy (anonymous):

make the answer interms of acceleration of the ball.. how hard denotes the force.. since mass of the ball is a constant.. let us assume it as ' m'.. this accelaration multiplied by ' m' gives the force

OpenStudy (anonymous):

the distances to be equated are .the point from where he starts running and the point from where the ball is thrown.. no other distances included

OpenStudy (anonymous):

becuase any distance travelled before throwing the ball does not affect the force

OpenStudy (anonymous):

how far the ball would be thrown? thats the displacement eqn for a projectile.. :)

OpenStudy (ashley1nonly):

i suppose because the player will only keep running until the ball stops.

OpenStudy (ashley1nonly):

when the ball stops he stops

OpenStudy (ashley1nonly):

i just figured since he ran a distance of D before the ball was thrown he also ran a distance of W to catch the ball that was thrown. so we would need D+W to calculate total distance.

OpenStudy (anonymous):

is he running D with the ball?

OpenStudy (anonymous):

what i meant was.. he is holding theball and running 5m forward... after 5m he throws the ball forward... lets assume the ball moves 5m forward.. so how many meters should he run?

OpenStudy (ashley1nonly):

No the ball is thrown when the player is D distance away but he still travels a short distance until he catches the ball.

OpenStudy (ashley1nonly):

the player throwing the ball is still

OpenStudy (anonymous):

there are 2 ppl?

OpenStudy (anonymous):

can you explain the scenario

OpenStudy (ashley1nonly):

There are two people playing football The quarter back throws the football when the runner back gets a distance of D.

OpenStudy (ashley1nonly):